# On Strongly Regular Graphs and the Friendship Theorem

### v1

@@
Abstract. This paper presents an alternative proof of the celebrated friendship theorem, originally established by Erdős, Rényi, and Sós (1966). The proof is based on a closed-form expression for the Lovász $\vartheta$-function of strongly regular graphs, which was recently derived by the author. Leveraging this expression, the paper also establishes new necessary conditions for a strongly regular graph to be a spanning subgraph or an induced subgraph of another strongly regular graph. In the case of induced subgraphs, the analysis further incorporates a property of graph energies.
Keywords. Friendship theorem, strongly regular graph, Lovász $\vartheta$-function, graph invariants, graph energy, windmill graph, spanning subgraph, induced subgraph.
@@ 1 Introduction
The friendship theorem in graph theory states that if a finite graph has the property that every pair of distinct vertices shares exactly one common neighbor, then the graph must be a collection of triangles that share a common vertex. In other words, the theorem asserts that such a finite graph is necessarily a friendship graph (also known as a windmill graph), meaning that it consists of a central vertex connected to all others with additional edges forming triangles. The name of this theorem stems from its human interpretation, where if every two individuals in a group of people have exactly one mutual friend, then there must be someone in that group who is everybody’s friend (often referred to as the ”politician”).
@@
That theorem was first proved by Erdős, Rényi, and Sós [5], and it found diverse applications in combinatorics and other disciplines. While the friendship theorem itself is a result in graph theory, its structural properties make it useful in designing graph-based codes, network coding, and cryptographic schemes. In network coding, graphs represent information flow between nodes, and the friendship property ensures that all transmissions go through a single relay node, which models centralized communication structures such as satellite communication, where every ground station interacts via a single hub. This property is also useful in wireless sensor networks, where multiple sensors communicate through a central processing node. Furthermore, friendship graphs can be linked to block designs and combinatorial structures used in constructing error-correcting codes with good covering properties. Beyond coding, the friendship theorem has a broad range of applications in networks, biology, computer science, social sciences, and economics. The fact that every two vertices share exactly one mutual neighbor makes it a useful model for systems where a single hub controls multiple independent interactions, such as in social influence, network protocols, and centralized coordination problems.
@@
The original proof of the theorem involves arguments based on graph connectivity, degree counting, and combinatorial reasoning. The reader is referred to a nice exposition of a proof of the friendship theorem in Chapter 44 of [1], which combines combinatorics and linear algebra (spectral graph theory). Due to its importance and simplicity, the theorem has been proved in various ways. The reader is also referred to [13] for historical notes on its proofs, as well as a new combinatorial proof proposed therein.
@@
The aim of this paper is to provide a different proof of that fundamental theorem, which relies on a recent finding of the author [14], where a closed-form expression of the celebrated Lovász $\vartheta$-function has been derived for the structured family of strongly regular graphs. We believe that this proof offers a new perspective to the available collection of proofs of the friendship theorem. Leveraging this expression, the paper also establishes new necessary conditions for a strongly regular graph to be a spanning subgraph or an induced subgraph of another strongly regular graph. In the case of induced subgraphs, the analysis further incorporates a property of graph energies. In general, understanding whether a graph is a spanning subgraph or an induced subgraph of another given graph is a meaningful and widely applicable question across theoretical, algorithmic, and applied fields of graph theory.
@@
This paper is structured as follows. Section 2 provides some essential preliminaries for our alternative proof of the friendship theorem in Section 3. A discussion on a spectral graph-theoretic proof that is closely related to the proof presented in [1, Chapter 44] is also provided in Section 3. Based on the Lovász $\vartheta$-function of strongly regular graphs, Sections 4 and 5 establish new necessary conditions for one strongly regular graph to be either a spanning subgraph or an induced subgraph, respectively, of another strongly regular graph. The proposed conditions can be easily checked, and their utility is exemplified.
@@ 2 Preliminaries
where the minimum is taken over all orthonormal representations $\{{\bf{u}}_{i}:i\in\operatorname{\mathsf{V}}(\mathsf{G})\}$ of $\mathsf{G}$, and over all unit vectors ${\bf{c}}$. The unit vector $\bf{c}$ is called the handle of the orthonormal representation. By the Cauchy-Schwarz inequality,
$\bigl{|}{\bf{c}}^{\mathrm{T}}{\bf{u}}_{i}\bigr{|}\leq\|{\bf{c}}\|\,\|{\bf{u}}_{i}\|=1\;\Longrightarrow\;\vartheta(\mathsf{G})\geq 1,$
with equality if and only if $\mathsf{G}$ is a complete graph.
Without loss of generality, the dimension $d$ can be set to be equal to the order of $\mathsf{G}$, i.e., $d=|\operatorname{\mathsf{V}}(\mathsf{G})|$.
Let
  • ${\bf{A}}$ is the $n\times n$ adjacency matrix of $\mathsf{G}$   ($n\triangleq|\operatorname{\mathsf{V}}(\mathsf{G})|$);
@@
  • ${\bf{\mathcal{S}}}_{+}^{n}$ is the set of all $n\times n$ positive semidefinite matrices.
The following semidefinite program (SDP) can be used for computing $\vartheta(\mathsf{G})$ (see [11] and [12, Chapter 11]):
@@
This renders the computational complexity of $\vartheta(\mathsf{G})$ feasible. More explicitly, there exist some standard algorithms in convex optimization that numerically compute $\vartheta(\mathsf{G})$, for every graph $\mathsf{G}$, with precision of $r$ decimal digits, and in polynomial-time in $n$ and $r$.
Let $\alpha(\mathsf{G})$, $\operatorname{\omega}(\mathsf{G})$, and $\operatorname{\chi}(\mathsf{G})$ denote the independence number, clique number, and chromatic number of a graph $\mathsf{G}$. Then, the Lovász $\vartheta$-function of $\mathsf{G}$ and its complement graph $\overline{\mathsf{G}}$ satisfy the following properties (see [11, 12]):
@@
    $\alpha(\mathsf{G})\leq\vartheta(\mathsf{G})\leq\operatorname{\chi}(\overline{\mathsf{G}}),$ (3)
    $\operatorname{\omega}(\mathsf{G})\leq\vartheta(\overline{\mathsf{G}})\leq\operatorname{\chi}(\mathsf{G}).$ (4)
@@
    Computational complexity: The graph invariants $\alpha(\mathsf{G})$, $\operatorname{\omega}(\mathsf{G})$, and $\operatorname{\chi}(\mathsf{G})$ are NP-hard problems. However, as mentioned above, the numerical computation of $\vartheta(\mathsf{G})$ is in general feasible.
@@
Let $\mathsf{G}$ be a $d$-regular graph of order $n$. The graph $\mathsf{G}$ is a strongly regular graph (SRG) if there also exist nonnegative integers $\lambda$ and $\mu$ such that
  • Every pair of adjacent vertices have exactly $\lambda$ common neighbors;
  • Every pair of distinct and non-adjacent vertices have exactly $\mu$ common neighbors.
Such a strongly regular graph is said to be a graph within the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$.
It should be noted that within the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$, with fixed four parameters $n$, $d$, $\lambda$, and $\mu$, there may exist several strongly regular graphs that are nonisomorphic to each other. For example, according to [3], there exist 167 nonisomorphic strongly regular graphs within the family $\operatorname{\mathsf{srg}}(64,18,2,6)$.
@@
As a common sufficient condition, note that all the inequalities in (6) and (7) hold with equality if $\mathsf{G}$ is a strongly regular graph. (Recall that a graph $\mathsf{G}$ is strongly regular if and only if its complement $\overline{\mathsf{G}}$ is so). The following result provides a closed-form expression of the Lovász $\theta$-function of all strongly regular graphs.
Corollary 2.1 (The Lovász $\vartheta$-Function of Strongly Regular Graphs, [14]).
Let $\mathsf{G}$ be a strongly regular graph within the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$. Then,
@@
In light of Corollary 2.1, all strongly regular graphs within a family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ are not only cospectral graphs [2] but they also share the same value of the Lovász $\vartheta$-function. In that context, it should be noted that there exist connected, cospectral, and irregular nonisomorphic graphs with an identical vertex set that are jointly cospectral with respect to several of their associated matrices (including the adjacency matrix) and share identical values of certain graph invariants, such as the independence number, clique number, and chromatic number, and yet they differ in their Lovász $\vartheta$-functions. The reader is referred to [15] for such explicit graph constructions and related analysis.
Corollary 2.2 (A Relation for Strongly Regular Graphs, [14]).
The relation
$\vartheta(\mathsf{G})\,\vartheta(\overline{\mathsf{G}})=n,$ (11)
holds not only for all vertex-transitive graphs (see [11, Theorem 8]), but it also holds for all strongly regular graphs (in general, strongly regular graphs may not be vertex-transitive, see, e.g., [6]).
3 The Friendship Theorem: An Alternative Proof
This paper presents an alternative proof of the celebrated friendship theorem, originally established in [5], which relies on the Lovász $\vartheta$-function of strongly regular graphs in Corollary 2.1.

### v3

@@
Abstract. This paper presents an alternative proof of the celebrated friendship theorem, originally established by Erdős, Rényi, and Sós (1966). The proof relies on a closed-form expression for the Lovász $\vartheta$-function of strongly regular graphs, recently derived by the author. Additionally, the paper considers some known extensions of the theorem, offering discussions that provide insights into the friendship theorem, one of its extensions, and the proposed proof. Leveraging the closed-form expression for the Lovász $\vartheta$-function of strongly regular graphs, the paper further establishes new necessary conditions for a strongly regular graph to be a spanning or induced subgraph of another strongly regular graph. In the case of induced subgraphs, the analysis also incorporates a property of graph energies. Some of these results are extended to regular graphs and their subgraphs.
Keywords. Friendship theorem, strongly regular graph, Lovász $\vartheta$-function, windmill graph, graph invariants, graph energy, spanning subgraph, induced subgraph.
@@ 1 Introduction
The friendship theorem in graph theory states that if a finite graph has the property that every pair of distinct vertices shares exactly one common neighbor, then the graph consists of edge-disjoint triangles sharing a single, central vertex. In other words, such a graph is necessarily a friendship graph (also called a windmill graph), where a central vertex is adjacent to all others, and additional edges form edge-disjoint triangles. The theorem’s name reflects its intuitive human interpretation, which asserts that if every two individuals in a (finite) group have exactly one mutual friend, then there must be someone who is everybody’s friend (often referred to as the ”politician”).
@@
That theorem was first proved by Erdős, Rényi, and Sós [1], and it found diverse applications in combinatorics and other disciplines. While the friendship theorem itself is a result in graph theory, its structural properties make it useful in designing graph-based codes, network coding, and cryptographic schemes. In network coding, graphs represent information flow between nodes, and the friendship property ensures that all transmissions go through a single relay node, which models centralized communication structures such as satellite communication, where every ground station interacts via a single hub. This property is also useful in wireless sensor networks, where multiple sensors communicate through a central processing node. Furthermore, friendship graphs can be linked to block designs and combinatorial structures used to construct error-correcting codes with good covering properties. Beyond these applications, the friendship theorem has a broad range of applications in computer science, biology, social sciences, and economics. The fact that every two vertices share exactly one mutual neighbor makes it a useful model for systems where a single hub controls multiple independent interactions, such as in social influence, network protocols, and centralized coordination problems.
@@
The original proof of the theorem involves arguments based on graph connectivity, degree counting, and combinatorial reasoning. The reader is referred to a nice exposition of a proof of the friendship theorem in Chapter 44 of [2], which combines combinatorics and linear algebra (spectral graph theory). Due to its importance and simplicity, the theorem has been proved in various ways. For historical insights into these proofs, as well as a novel combinatorial proof and an extension of the theorem, the reader is referred to [3].
@@
The aim of this paper is twofold. First, it provides an alternative proof of that theorem, relying on a recent result by the author [4], in which a closed-form expression for the celebrated Lovász $\vartheta$-function was derived for the structured family of strongly regular graphs. We believe this proof offers a new perspective on the existing collection of proofs of the friendship theorem. The paper further considers some known extensions of the theorem, offering discussions that provide insights into the friendship theorem, one of its extensions from [3], and the proposed proof. The second aim of this paper, building on the closed-form expression for the Lovász $\vartheta$-function of all strongly regular graphs, is to establish new necessary conditions for a strongly regular graph to be a spanning or induced subgraph of another strongly regular graph. For induced subgraphs, the analysis also incorporates a property of graph energies. Some of these results are extended to regular graphs and their subgraphs. In general, determining whether a graph is a spanning or induced subgraph of another given graph is an important and broadly applicable problem across theoretical, algorithmic, and applied areas of graph theory.
@@
This paper is structured as follows. Section 2 introduces the notation and essential background required for the analysis in this paper. Section 3 provides our alternative proof of the friendship theorem, followed by a variation of a known spectral graph-theoretic proof, and a consideration of the friendship theorem and our alternative proof to gain further insights into the theorem, our proposed proof, and one of the theorem’s extensions in [3]. Based on the Lovász $\vartheta$-function of strongly regular graphs, Sections 4 and 5 establish new necessary conditions for one strongly regular graph to be a spanning or induced subgraph, respectively, of another strongly regular graph. The proposed conditions can be easily checked, their utility is demonstrated, and they are extended to regular graphs and subgraphs, with a consideration of their computational complexity.
@@ 2 Preliminaries
This section presents the notation and necessary background, including definitions and theorems essential for the analysis in this paper.
Let $\mathsf{G}$ be a graph with vertex set $\operatorname{\mathsf{V}}(\mathsf{G})$ and edge set $\operatorname{\mathsf{E}}(\mathsf{G})$. A graph $\mathsf{G}$ is called simple if it has no self-loops and no multiple edges between any pair of vertices. Throughout the paper, unless explicitly mentioned, all graphs under consideration are simple, finite, and undirected. The following standard notation and definitions are used.
Definition 2.1 (Graph complement).
The complement of a graph $\mathsf{G}$, denoted by $\overline{\mathsf{G}}$, is a graph whose vertex set is $\operatorname{\mathsf{V}}(\mathsf{G})$, and its edge set is the complement set $\overline{\mathsf{\operatorname{\mathsf{E}}(\mathsf{G})}}$. Every vertex in $\operatorname{\mathsf{V}}(\mathsf{G})$ is nonadjacent to itself in $\mathsf{G}$ and $\overline{\mathsf{G}}$, so $\{i,j\}\in\operatorname{\mathsf{E}}(\overline{\mathsf{G}})$ if and only if $\{i,j\}\notin\operatorname{\mathsf{E}}(\mathsf{G})$ with $i\neq j$.
Definition 2.2 (Integer-valued graph invariants).
  • Let $k\in\mathbb{N}$. A proper $k$-coloring of a graph $\mathsf{G}$ is a function $c\colon\operatorname{\mathsf{V}}(\mathsf{G})\to\{1,2,...,k\}$, where $c(v)\neq c(u)$ for every $\{u,v\}\in\operatorname{\mathsf{E}}(\mathsf{G})$. The chromatic number of $\mathsf{G}$, denoted by $\chi(\mathsf{G})$, is the smallest $k$ for which there exists a proper $k$-coloring of $\mathsf{G}$.
  • A clique in a graph $\mathsf{G}$ is a subset of vertices $U\subseteq\operatorname{\mathsf{V}}(\mathsf{G})$, where $\{u,v\}\in\operatorname{\mathsf{E}}(\mathsf{G})$ for every $u,v\in U$ with $u\neq v$. The clique number of $\mathsf{G}$, denoted by $\omega(\mathsf{G})$, is the largest size of a clique in $\mathsf{G}$.
  • An independent set in a graph $\mathsf{G}$ is a subset of vertices $U\subseteq\operatorname{\mathsf{V}}(\mathsf{G})$, where $\{u,v\}\notin\operatorname{\mathsf{E}}(\mathsf{G})$ for every $u,v\in U$. The independence number of $\mathsf{G}$, denoted by $\alpha(\mathsf{G})$, is the largest size of an independent set in $\mathsf{G}$. Consequently, $\alpha(\mathsf{G})=\omega(\overline{\mathsf{G}})$ for every graph $\mathsf{G}$.
These integer-valued functions of a graph are invariant under graph isomorphisms, so they are referred to as graph invariants.
We next introduce the Lovász $\vartheta$-function of a graph $\mathsf{G}$, and then consider some of its useful properties. To that end, orthonormal representations of graphs are introduced.
Definition 2.3 (Orthonormal representations of a graph).
Let $\mathsf{G}$ be a finite, simple, and undirected graph, and $d\in\mathbb{N}$. An orthonormal representation of $\mathsf{G}$ in the $d$-dimensional Euclidean space $\mathbb{R}^{d}$ assigns to each vertex $i\in\operatorname{\mathsf{V}}(\mathsf{G})$ a unit vector ${\bf{u}}_{i}\in\mathbb{R}^{d}$ such that, for every two distinct and nonadjacent vertices in $\mathsf{G}$, their assigned vectors are orthogonal.
@@
where the minimum is taken over all orthonormal representations $\{{\bf{u}}_{i}:i\in\operatorname{\mathsf{V}}(\mathsf{G})\}$ of $\mathsf{G}$, and over all unit vectors ${\bf{c}}$. The unit vector $\bf{c}$ is called the handle of the orthonormal representation. By the Cauchy-Schwarz inequality, $\bigl{|}{\bf{c}}^{\mathrm{T}}{\bf{u}}_{i}\bigr{|}\leq\|{\bf{c}}\|\,\|{\bf{u}}_{i}\|=1$, so $\vartheta(\mathsf{G})\geq 1$, with equality if and only if $\mathsf{G}$ is a complete graph.
The dimension $d$ of the Euclidean space $\mathbb{R}^{d}$ in the orthonormal representations of $\mathsf{G}$, over which the minimization on the right-hand side of (1) is performed can be set to the order of $\mathsf{G}$, i.e., $d=|\operatorname{\mathsf{V}}(\mathsf{G})|$ (see p. 183 of [6]). Let the following notation be used:
  • ${\bf{A}}$ is the $n\times n$ adjacency matrix of $\mathsf{G}$   ($n\triangleq|\operatorname{\mathsf{V}}(\mathsf{G})|$), where $A_{i,j}=1$ if and only if $\{i,j\}\in\operatorname{\mathsf{E}}(\mathsf{G})$ and $A_{i,j}=0$ otherwise;
@@
  • ${\bf{\mathcal{S}}}_{+}^{n}$ is the set of all $n\times n$ positive semidefinite matrices;
  • The eigenvalues of ${\bf{A}}$ are given in decreasing order by
    $\lambda_{\max}(\mathsf{G})=\lambda_{1}(\mathsf{G})\geq\lambda_{2}(\mathsf{G})\geq\ldots\geq\lambda_{n}(\mathsf{G})=\lambda_{\min}(\mathsf{G}).$ (2)
  • The adjacency spectrum of $\mathsf{G}$ is the multiset of the eigenvalues of ${\bf{A}}$, counted with multiplicities.
The following semidefinite program (SDP) computes $\vartheta(\mathsf{G})$ (by Theorem 4 of [5]):
@@
This renders the computational complexity of $\vartheta(\mathsf{G})$ feasible. Specifically, there exist standard algorithms in convex optimization that numerically compute $\vartheta(\mathsf{G})$, for every graph $\mathsf{G}$, with precision of $r$ decimal digits, and in polynomial-time in $n$ and $r$ (see Section 11.3 of [6]).
The Lovász $\vartheta$-function of $\mathsf{G}$ and its complement graph $\overline{\mathsf{G}}$ satisfy the following [5, 6]:
@@
    $\alpha(\mathsf{G})\leq\vartheta(\mathsf{G})\leq\chi(\overline{\mathsf{G}}),$ (4)
    $\omega(\mathsf{G})\leq\vartheta(\overline{\mathsf{G}})\leq\chi(\mathsf{G}).$ (5)
@@
    Computational complexity: The graph invariants $\alpha(\mathsf{G})$, $\omega(\mathsf{G})$, and $\chi(\mathsf{G})$ are NP-hard problems. However, as mentioned above, the numerical computation of $\vartheta(\mathsf{G})$ is in general feasible, so it provides polynomial-time computable bounds on very useful graph invariants that are hard to compute.
@@
Let $\mathsf{G}$ be a $d$-regular graph of order $n$. The graph $\mathsf{G}$ is a strongly regular graph (SRG) if there exist nonnegative integers $\lambda$ and $\mu$ such that
  • Every pair of adjacent vertices has exactly $\lambda$ common neighbors;
  • Every pair of distinct, non-adjacent vertices has exactly $\mu$ common neighbors.
Such a strongly regular graph is said to be a graph in the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$.
It is important to note that these four parameters are interrelated (see Proposition 3.1 for known results). Consequently, for some parameter vectors $(n,d,\lambda,\mu)$, the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ contains no graphs. Furthermore, for parameter vectors $(n,d,\lambda,\mu)$ such that the set $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ is nonempty, there may exist several nonisomorphic strongly regular graphs within it. For example, according to [7], there exist 167 nonisomorphic strongly regular graphs in the family $\operatorname{\mathsf{srg}}(64,18,2,6)$. The reader is referred to [7, 8, 9] for the enumeration of strongly regular graphs with given parameter vectors.
Proposition 2.1.
[9] A graph $\mathsf{G}$ is strongly regular if and only if its complement $\overline{\mathsf{G}}$ is so. Furthermore, if $\mathsf{G}$ is a strongly regular graph in the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$, then $\overline{\mathsf{G}}$ is a strongly regular graph in the family $\operatorname{\mathsf{srg}}(n,n-d-1,n-2d+\mu-2,n-2d+\lambda)$.
@@
As a common sufficient condition, note that all the inequalities in (7) and (8) hold with equality if $\mathsf{G}$ is a strongly regular graph. The following result provides a closed-form expression of the Lovász $\vartheta$-function of all strongly regular graphs.
Theorem 2.2 (The Lovász $\vartheta$-function of strongly regular graphs, [4]).
Let $\mathsf{G}$ be a strongly regular graph in the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$. Then,
@@
In light of Theorem 2.2, all strongly regular graphs in a family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ are not only cospectral [9], but they also share the same value of the Lovász $\vartheta$-function.
Corollary 2.1 ([4]).
Let $\mathsf{G}$ be a strongly regular graph on $n$ vertices. Then,
$\vartheta(\mathsf{G})\,\vartheta(\overline{\mathsf{G}})=n.$ (12)
Remark 2.2.
Corollary 2.1 shows that equality (12) does not only for all vertex-transitive graphs (see Theorem 8 of [5]), but it also holds for all strongly regular graphs. It should be noted that strongly regular graphs may not be vertex-transitive (see, e.g., [10]).
The closed-form expression in Theorem 2.2 for the Lovász $\vartheta$-function of strongly regular graphs eliminates the need for a numerical solution to the SDP in (3), while also providing an explicit analytical expression for that function. This facilitates the derivation of analytical bounds on key graph invariants, which are generally NP-hard to compute. These bounds are expressed in terms of the parameters $(n,d,\lambda,\mu)$ characterizing strongly regular graphs, as it is next shown in Corollary 2.2.
Corollary 2.2 (Bounds on graph invariants of strongly regular graphs, [4]).
Let $\mathsf{G}$ be a strongly regular graph in the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$. Then,
$\alpha(\mathsf{G})\leq\bigg{\lfloor}\dfrac{n\,(t+\mu-\lambda)}{2d+t+\mu-\lambda}\bigg{\rfloor}$ (13)
$\omega(\mathsf{G})\leq 1+\bigg{\lfloor}\dfrac{2d}{t+\mu-\lambda}\bigg{\rfloor},$ (14)
$\chi(\mathsf{G})\geq 1+\bigg{\lceil}\dfrac{2d}{t+\mu-\lambda}\bigg{\rceil},$ (15)
$\chi(\overline{\mathsf{G}})\geq\bigg{\lceil}\dfrac{n\,(t+\mu-\lambda)}{2d+t+\mu-\lambda}\bigg{\rceil},$ (16)
where $t$ is given in (11).
Proof.
The bounds in (13)–(16) follow from the combination of the sandwich theorem in (4) and (5) with Theorem 2.2. ∎
Example 2.1 (Bounds on graph invariants of strongly regular graphs).
The tightness of the bounds in Corollary 2.2 is exemplified for four strongly regular graphs as follows.
  1. (1)
    Petersen graph: Let $\mathsf{G}_{1}$ be the Petersen graph, the unique strongly regular graph in the family $\operatorname{\mathsf{srg}}(10,3,0,1)$ (see Section 10.3 of [9]). For $\mathsf{G}_{1}$, the upper bounds on its independence and clique numbers in (13) and (14), respectively, as well as the lower bound on its chromatic number in (15) are tight:
    $\alpha(\mathsf{G}_{1})=4,\quad\omega(\mathsf{G}_{1})=2,\quad\chi(\mathsf{G}_{1})=3.$ (17)
  2. (2)
    Schläfli, Shrikhande, and Hall-Janko graphs:
    • The Schläfli graph is (up to isomorphism) the unique strongly regular graph in the family $\operatorname{\mathsf{srg}}(27,16,10,8)$ (see Section 10.7 of [9]).
    • The Shrikhande graph is one of two nonisomorphic strongly regular graphs in the family $\operatorname{\mathsf{srg}}(16,6,2,2)$ (see Section 10.6 of [9]).
    • The Hall-Janko graph is the unique strongly regular graph in the family $\operatorname{\mathsf{srg}}(100,36,14,12)$ (see Section 10.32 of [9]).
    Let $\mathsf{G}_{2}$, $\mathsf{G}_{3}$, and $\mathsf{G}_{4}$ denote these graphs, respectively. The lower bounds on the chromatic numbers of these graphs, as given in (15), are all tight:
    $\chi(\mathsf{G}_{2})=9,\quad\chi(\mathsf{G}_{3})=4,\quad\chi(\mathsf{G}_{4})=10.$ (18)
  3. (3)
    For the Shrikhande graph ($\mathsf{G}_{3}$),
    • the upper bound on its independence number in (13) is tight: $\alpha(\mathsf{G}_{3})=4$,
    • the upper bound on its clique number in (14) is, however, not tight as the upper bound is 4, while the actual clique number is $\omega(\mathsf{G}_{3})=3$.
3 The Friendship Theorem: An Alternative Proof and Further Extensions
This section is structured as follows: it presents the renowned friendship theorem, originally established in [1] (see Section 3.1), and it provides a new alternative proof that relies on the Lovász $\vartheta$-function of strongly regular graphs (see Theorem 2.2 and Section 3.2). This section further provides a variation of a proof from Chapter 44 of [2], relying on spectral properties of strongly regular graphs (see Section 3.3). It finally considers further generalizations, providing discussions in order to get further insights into the theorem, our proposed proof, and one of the theorem’s extensions in [3], which are also supported by numerical results for (small and large) strongly regular graphs (see Section 3.4).
3.1 The friendship theorem